# Lightweight Dataset Pruning without Full Training via Example Difficulty and Prediction Uncertainty

### v1

@@ Abstract
Recent advances in deep learning rely heavily on massive datasets, leading to substantial storage and training costs. Dataset pruning aims to alleviate this demand by discarding redundant examples. However, many existing methods require training a model with a full dataset over a large number of epochs before being able to prune the dataset, which ironically makes the pruning process more expensive than just training the model on the entire dataset. To overcome this limitation, we introduce a Difficulty and Uncertainty-Aware Lightweight (DUAL) score, which aims to identify important samples from the early training stage by considering both example difficulty and prediction uncertainty. To address a catastrophic accuracy drop at an extreme pruning, we further propose a ratio-adaptive sampling using Beta distribution. Experiments on various datasets and learning scenarios such as image classification with label noise and image corruption, and model architecture generalization demonstrate the superiority of our method over previous state-of-the-art (SOTA) approaches. Specifically, on ImageNet-1k, our method reduces the time cost for pruning to 66% compared to previous methods while achieving a SOTA, specifically 60% test accuracy at a 90% pruning ratio. On CIFAR datasets, the time cost is reduced to just 15% while maintaining SOTA performance.  [^1] Our codebase is available at github.com/behaapyy/dual-pruning .
@@ 1 Introduction
However, many existing pruning methods require training a model with a full dataset over a number of epochs to measure the importance of each sample, which ironically makes the pruning process more expensive than just training the model once on the original large dataset. For instance, several score-based methods (Toneva et al., 2018; Pleiss et al., 2020; Gordon et al., 2021; He et al., 2024; Zhang et al., 2024) require training as they utilize the dynamics from the whole training process. Some geometry-based methods, (Xia et al., 2022; Yang et al., 2024) leverage features from the penultimate layer of the trained model, therefore training a model is also required. Hybrid methods (Zheng et al., 2022; Maharana et al., 2023; Tan et al., 2025), which address the difficulty and diversity of samples simultaneously, still hold the same limitation as they use existing score metrics. Having to compute the dot product of learned features to get the neighborhood information makes them even more expensive to utilize.
@@ 2 Related Works
Score-based. EL2N (Gordon et al., 2021) calculates $L2$ norms of the error vector as an approximation of the gradient norm. Entropy (Coleman et al., 2020) quantifies the information contained in the predicted probabilities at the end of training. However, the outcomes of such “snapshot” methods differ significantly from run to run, making it difficult to obtain a reliable score in a single run, as can be seen in Figure 7, Appendix B.
@@ 3.2 Difficulty & Uncertainty-Aware Lightweight Score
A theoretical analysis of a toy example further verifies the intuition above. Consider a linearly separable binary classification task $\left\{({\bm{x}}_{i}\in{\mathbb{R}}^{n},y_{i}\in\{\pm 1\})\right\}_{i=1}^{N}$, where $N=2$ with $\lVert{\bm{x}}_{1}\rVert$ $\ll$ $\langle{\bm{x}}_{1},{\bm{x}}_{2}\rangle<\lVert{\bm{x}}_{2}\rVert$. Without loss of generality, we set $y_{1}=y_{2}=+1$. A linear classifier, $f({\bm{x}};{\bm{w}})={\bm{w}}^{\top}{\bm{x}}$, is employed as the model in our analysis. The parameter ${\bm{w}}$ is initialized at zero and updated by gradient descent. Soudry et al. (2018) prove that the parameter of linear classifiers diverges to infinity, but directionally converges to the $L_{2}$ maximum margin separator. This separator is determined by the support vectors closest to the decision boundary. If a valid pruning method encounters this task, then it should retain the point closer to the decision boundary, which is ${\bm{x}}_{1}$ in our case, and prune ${\bm{x}}_{2}$.
@@ 4 Experiments
Pruning Rate ($\rightarrow$) 30% 50% 70% 80% 90% 30% 50% 70% 80% 90%
@@ 4.1 Experimental Settings
Baselines. The baselines considered in this study are listed as follows [^2] Infomax (Tan et al., 2025) was excluded as it employs different base hyperparameters in the original paper compared to other baselines and does not provide publicly available code. See Appendix A.1 for more discussion.: (1) Random; (2) Entropy (Coleman et al., 2020); (3) Forgetting (Toneva et al., 2018); (4) EL2N (Gordon et al., 2021); (5) AUM (Pleiss et al., 2020); (6) Moderate (Xia et al., 2022); (7) Dyn-Unc (He et al., 2024); (8) TDDS (Zhang et al., 2024); (9) CCS (Zheng et al., 2022); and (10) $\mathbb{D}^{2}$ (Maharana et al., 2023). To ensure a fair comparison, all methods were trained with the same base hyperparameters for training, and the best hyperparameters reported in their respective original works for scoring. Technical details are provided in the Section A.1.
@@ 4.3.1 Label Noise and Image Corruption
Data affected by label noise or image corruption are difficult and unnecessary samples that hinder model learning and degrade generalization performance. Therefore, filtering out these samples through data pruning is crucial. Most data pruning methods, however, either focus solely on selecting difficult samples based on example difficulty (Gordon et al., 2021; Pleiss et al., 2020; Coleman et al., 2020) or prioritize dataset diversity (Zheng et al., 2022; Xia et al., 2022), making them unsuitable for effectively pruning such noisy and corrupted samples.
@@ A.1 Details on Baseline Implementation
EL2N (Paul et al., 2021) is defined as the error $L2$ norm between the true labels and predictions of the model. The examples with low scores are pruned out. We calculated error norms at epoch 20 from five independent runs, then the average was used for Ethe L2N score.
@@ B.1 Image Classification with Label Noise
We evaluated the performance of our proposed method across a wide range of pruning levels, from 10% to 90%, and compared the final accuracy with that of baseline methods. As shown in the Table 5-8, our method consistently outperforms the competition with a substantial margin in most cases. For a comprehensive analysis of performance under noisy conditions, please refer to Tables 5 to 7 for CIFAR-100, which show results for 20%, 30%, and 40% noise, respectively. Additionally, the results for 20% label noise in Tiny-ImageNet are shown in Table 8.
@@ B.2 Image Classification with Image Corruption
We also evaluated the robustness of our proposed method against five different types of realistic image corruption: motion blur, fog, reduced resolution, rectangular occlusion, and Gaussian noise across the corruption rate from 20% to 40%. The ratio of each type of corruption is 4% for 20% corruption, 6% for 30% corruption, and 8% for 40% corruption. Example images for each type of corruption can be found in Figure 12. Motion blur, reduced resolution, and rectangular occlusion are somewhat distinguishable, whereas fog and Gaussian noise are difficult for the human eye to differentiate. Somewhat surprisingly, our DUAL pruning prioritizes removing the most challenging examples, such as fog and Gaussian corrupted images, as shown in Figure 14.
@@ B.3 Cross-architecture generalization
In this section, we investigate the cross-architecture generalization ability of our proposed method. Specifically, we calculate the example score on one architecture and test its coreset performance on a different architecture. This evaluation aims to assess the ability of example scores to be transferred across diverse architectural designs.
@@ B.4 Effectiveness of Beta Sampling
Appendix C Detailed Explanation about Our Method
In this section, we provide details on the implementation used across all experiments for reproducibility. Appendix C.2 presents the full algorithm for our pruning method, DUAL, along with the Beta sampling strategy. Additionally, in a later subsection, we visualize the selected data using Beta sampling.
Recall that we define our sampling distribution $\mathrm{Beta}(\alpha_{r},\beta_{r})$ as follows:
$\begin{split}\beta_{r}&=C\left(1-\mu_{\mathcal{D}}\right)\left(1-r^{c_{\mathcal{D}}}\right)\\ \alpha_{r}&=C-\beta_{r},\end{split}$ (6)
where $\mu_{\mathcal{D}}\in[0,1]$ is the probability mean of the highest DUAL score training sample. To ensure stability, we compute this as the average probability mean of the 10 highest DUAL score training samples. Additionally, as mentioned earlier, we set the value of $C$ to 15 across all experiments. For technical details, we add 1 to $\alpha_{r}$ to further ensure that the PDF remains stationary at low pruning ratios.
We illustrate the Beta PDF, as defined above, in Figure 15 for different values of $c_{\mathcal{D}}$. In both subplots, we set $\mu_{\mathcal{D}}$ as 0.25. The left subplot shows the PDF with $c_{\mathcal{D}}=5.5$, which corresponds to the value used in CIFAR-10 experiments, while the right subplot visualizes the PDF where $c_{\mathcal{D}}=4$, corresponding to CIFAR-100.
Figure 15: Visualization of Beta distribution for varying $c_{\mathcal{D}}$. The left subplot corresponds to the value used in CIFAR-10, and the right subplot corresponds to the value used in CIFAR-100.
@@ D.2 Experimental Results under Synthetic Setting
This section displays the figures plotted from the experiments on the synthetic dataset. We choose ${\mathcal{X}}={\mathbb{R}}^{2}$ and ${\mathcal{D}}=\left\{\ ((0.1,0.1),1),((10,5),1)\right\}$. We fix $J=10$ and $\eta=0.01$ (unless specified). The total time of training $T$ is specified for each figure for neat visualization. In this setting, the upper bound for the learning rate is $\log(75)/126.5\approx 0.034$.

### v2

@@ Abstract
Recent advances in deep learning rely heavily on massive datasets, leading to substantial storage and training costs. Dataset pruning aims to alleviate this demand by discarding redundant examples. However, many existing methods require training a model with a full dataset over a large number of epochs before being able to prune the dataset, which ironically makes the pruning process more expensive than just training the model on the entire dataset. To overcome this limitation, we introduce a Difficulty and Uncertainty-Aware Lightweight (DUAL) score, which aims to identify important samples from the early training stage by considering both example difficulty and prediction uncertainty. To address a catastrophic accuracy drop at an extreme pruning, we further propose a ratio-adaptive sampling using Beta distribution. Experiments on various datasets and learning scenarios such as image classification with label noise and image corruption, and model architecture generalization demonstrate the superiority of our method over previous state-of-the-art (SOTA) approaches. Specifically, on ImageNet-1k, our method reduces the time cost for pruning to 66% compared to previous methods while achieving a SOTA, specifically 60% test accuracy at a 90% pruning ratio. On CIFAR datasets, the time cost is reduced to just 15% while maintaining SOTA performance.  [^1] Our codebase is available at github.com/dual-pruning.
@@ 1 Introduction
However, many existing pruning methods require training a model with a full dataset over a number of epochs to measure the importance of each sample, which ironically makes the pruning process more expensive than just training the model once on the original large dataset. For instance, several score-based methods (Toneva et al., 2018; Pleiss et al., 2020; Paul et al., 2021; He et al., 2024; Zhang et al., 2024) require training as they utilize the dynamics from the whole training process. Some geometry-based methods, (Xia et al., 2022; Yang et al., 2024) leverage features from the penultimate layer of the trained model, therefore training a model is also required. Hybrid methods (Zheng et al., 2022; Maharana et al., 2023; Tan et al., 2025), which address the difficulty and diversity of samples simultaneously, still hold the same limitation as they use existing score metrics. Having to compute the dot product of learned features to get the neighborhood information makes them even more expensive to utilize.
@@ 2 Related Works
Score-based. EL2N (Paul et al., 2021) calculates $L2$ norms of the error vector as an approximation of the gradient norm. Entropy (Coleman et al., 2020) quantifies the information contained in the predicted probabilities at the end of training. However, the outcomes of such “snapshot” methods differ significantly from run to run, making it difficult to obtain a reliable score in a single run, as can be seen in Figure 7, Appendix B.
@@ 3.2 Difficulty & Uncertainty-Aware Lightweight Score
A theoretical analysis of a toy example further verifies the intuition above. Consider a linearly separable binary classification task $\left\{({\bm{x}}_{i}\in{\mathbb{R}}^{n},y_{i}\in\{\pm 1\})\right\}_{i=1}^{N}$, where $N=2$ with $\lVert{\bm{x}}_{1}\rVert^{2}$ $\ll$ $\langle{\bm{x}}_{1},{\bm{x}}_{2}\rangle<\lVert{\bm{x}}_{2}\rVert^{2}$. Without loss of generality, we set $y_{1}=y_{2}=+1$. A linear classifier, $f({\bm{x}};{\bm{w}})={\bm{w}}^{\top}{\bm{x}}$, is employed as the model in our analysis. The parameter ${\bm{w}}$ is initialized at zero and updated by gradient descent. Soudry et al. (2018) prove that the parameter of linear classifiers diverges to infinity, but directionally converges to the $L_{2}$ maximum margin separator. This separator is determined by the support vectors closest to the decision boundary. If a valid pruning method encounters this task, then it should retain the point closer to the decision boundary, which is ${\bm{x}}_{1}$ in our case, and prune ${\bm{x}}_{2}$.
@@ 4 Experiments
chat Pruning Rate ($\rightarrow$) 30% 50% 70% 80% 90% 30% 50% 70% 80% 90%
@@ 4.1 Experimental Settings
Baselines. The baselines considered in this study are listed as follows [^2] Infomax (Tan et al., 2025) was excluded as it employs different base hyperparameters in the original paper compared to other baselines and does not provide publicly available code. See Appendix A.1 for more discussion.: (1) Random; (2) Entropy (Coleman et al., 2020); (3) Forgetting (Toneva et al., 2018); (4) EL2N (Paul et al., 2021); (5) AUM (Pleiss et al., 2020); (6) Moderate (Xia et al., 2022); (7) Dyn-Unc (He et al., 2024); (8) TDDS (Zhang et al., 2024); (9) CCS (Zheng et al., 2022); and (10) $\mathbb{D}^{2}$ (Maharana et al., 2023). To ensure a fair comparison, all methods were trained with the same base hyperparameters for training, and the best hyperparameters reported in their respective original works for scoring. Technical details are provided in the Section A.1.
@@ 4.3.1 Label Noise and Image Corruption
Data affected by label noise or image corruption are difficult and unnecessary samples that hinder model learning and degrade generalization performance. Therefore, filtering out these samples through data pruning is crucial. Most data pruning methods, however, either focus solely on selecting difficult samples based on example difficulty (Paul et al., 2021; Pleiss et al., 2020; Coleman et al., 2020) or prioritize dataset diversity (Zheng et al., 2022; Xia et al., 2022), making them unsuitable for effectively pruning such noisy and corrupted samples.
@@ 4.3.2 Cross-Architecture Generalization
Table 3: Cross-architecture generalization performance on CIFAR-100 from ResNet-18 to ResNet-50. We report an average of five runs. ‘R50 $\rightarrow$ R50’ stands for score computation on ResNet-50, as a baseline. Test accuracy with full dataset is 80.1%.
@@ A.1 Details on Baseline Implementation
EL2N (Paul et al., 2021) is defined as the error $L2$ norm between the true labels and predictions of the model. The examples with low scores are pruned out. We calculated error norms at epoch 20 from five independent runs, then the average was used for the EL2N score.
@@ B.1 Image Classification with Label Noise
We evaluated the performance of our proposed method across a wide range of pruning levels, from 10% to 90%, and compared the final accuracy with that of baseline methods. As shown in the Table 5-8, our method consistently outperforms the competition with a substantial margin in most cases. For a comprehensive analysis of performance under noisy conditions, please refer to Tables 5 to 7 for CIFAR-100, which show results for 20%, 30%, and 40% noise, respectively. Additionally, the results for 20% label noise in Tiny-ImageNet are in Table 8.
@@ B.2 Image Classification with Image Corruption
We also evaluated the robustness of our proposed method against five different types of realistic image corruption: motion blur, fog, reduced resolution, rectangular occlusion, and Gaussian noise across the corruption rate from 20% to 40%. The ratio of each type of corruption is 4% for 20% corruption, 6% for 30% corruption, and 8% for 40% corruption. Example images for each type of corruption can be found in Figure 12. Motion blur, reduced resolution, and rectangular occlusion are somewhat distinguishable, whereas fog and Gaussian noise are difficult for the human eye to differentiate. Surprisingly, our DUAL pruning prioritizes removing the most challenging examples, such as fog and Gaussian corrupted images, as shown in Figure 14.
@@ B.3 Cross-architecture generalization
In this section, we investigate the cross-architecture generalization ability of our proposed method. We calculate the example score on one architecture and test its coreset performance on a different architecture. This evaluation aims to assess the ability of example scores to be transferred across diverse architectural designs.
@@ B.4 Effectiveness of Beta Sampling
Appendix C Detailed Explanation about Beta Sampling
Here, we provide details on our choice of Beta sampling. Appendix C.1 shows the visualization of the selected data using the Beta sampling. Appendix C.2 presents the full algorithm for our DUAL pruning method with the suggested Beta sampling strategy.
We begin by explaining why the Beta distribution is selected as the sampling distribution. The domain of Beta distribution is $[0,1]$, which naturally aligns with the range of prediction means. Moreover, the probability density function (PDF) of the Beta distribution can be shaped to decay at both tails, ensuring that samples with extreme scores are rarely selected. While other distributions, such as Gaussian, could also be considered, their support spans ${\mathbb{R}}$. As a result, they may assign non-negligible probability to values far outside the desired range unless their standard deviation is made extremely small.
We define our sampling distribution $\mathrm{Beta}(\alpha_{r},\beta_{r})$ as follows:
$\begin{split}\beta_{r}&=15\left(1-\mu_{\mathcal{D}}\right)\left(1-r^{c_{\mathcal{D}}}\right)\\ \alpha_{r}&=15-\beta_{r},\end{split}$ (6)
where $\mu_{\mathcal{D}}\in[0,1]$ is the probability mean of the highest DUAL score training sample. To ensure stability, we compute this as the average probability mean of the 10 highest DUAL score training samples. Additionally, as mentioned in Appendix A.2, we set the value of $C$ to 15 across all experiments. Recall that the variance of a Beta distribution with parameters is given by $\frac{\alpha\beta}{(\alpha+\beta)^{2}(\alpha+\beta+1)}$, so increasing $C$ leads to a lower variance, an effect illustrated in Figure 15. This enables a more concentrated sampling distribution in a specific region, improving the effectiveness of the sampling process by reducing unnecessary spread. In an implementation, we add 1 to $\alpha_{r}$ to provide a more targeted sampling.
Figure 15: Visualization of Beta distribution for varying $C$. Large $C$ enables more concentrated targeting.
Now we justify our choice of parameters $\alpha_{r}$ and $\beta_{r}$ in the Beta distribution. When the pruning ratio is set to zero, $\alpha_{r}$ and $\beta_{r}$ are configured so that the mean of the Beta distribution (which is $\frac{\alpha}{\alpha+\beta}$) matches the prediction mean of the highest-scoring sample. This allows the sampling to focus on high-score samples at low pruning ratios. To gradually include easier samples as the pruning ratio increases, we set parameters $\alpha_{r}$ and $\beta_{r}$ depending on the pruning ratio. While BOSS (Acharya et al., 2024) adjusts these parameters such that the mode of the Beta distribution (which is $\frac{\alpha-1}{\alpha+\beta-2}$ for $\alpha,\beta>1$) scales linearly with the pruning ratio $r$, we adopt a non-linear scaling by raising $r$ to the power of $c_{D}$. This results in a PDF that is remains almost stationary at low pruning ratios, but gradually shifts toward easier samples in a polynomial manner as the pruning ratio increases.
The hyperparameter $c_{D}$ is chosen based on the relative complexity of the dataset. We assumed that a larger dataset, which may have more samples per class, tends to be relatively easier. A higher value of $c_{D}$ results in smaller $\beta_{r}$, thereby increasing the mean and reducing the variance of the Beta distribution. For more difficult datasets, sampling easier examples becomes more important, which justifies using a larger $c_{D}$. Figure 16 illustrates the Beta PDF for different values of $c_{\mathcal{D}}$. In both subplots, we set $\mu_{\mathcal{D}}$ as 0.25. The left subplot shows the PDF with $c_{\mathcal{D}}=5.5$, which corresponds to the value used in our CIFAR-10 experiments, while the right subplot shows the case where $c_{\mathcal{D}}=4$, which is used in CIFAR-100.
Figure 16: Visualization of Beta distribution for varying $c_{\mathcal{D}}$. Left subplot corresponds to the value used in CIFAR-10, and the right subplot corresponds to the value used in CIFAR-100.
@@ D.2 Experimental Results under Synthetic Setting
This section displays the figures plotted from the experiments on the synthetic dataset. We choose ${\mathcal{X}}={\mathbb{R}}^{2}$ and ${\mathcal{D}}=\left\{\ ((0.1,0.1),1),((10,5),1)\right\}$. We fix $J=10$ and $\eta=0.01$ (unless specified). The total time of training $T$ is specified for each figure for neat visualization.